OpenStudy (anonymous):

In the game of golf, a score of 0 is called even par. Scores over par are represented by positive numbers and scores under par are represented by negative numbers. Over the course of three weeks, a golfer’s scores on a course are +2, –1, 0, –1, and +3. What is the combined score of these games? Is this not 3? It says its incorrect... what am I doing wrong

OpenStudy (anonymous):

i got 3 also but i would put the answer as +3

OpenStudy (mathstudent55):

+2 + (-1) + 0 + (-1) + 3 = 1 + 0 + (-1) + 3 = 1 + (-1) + 3 = 0 + 3 = 3

OpenStudy (anonymous):

Yea +3 makes sense.